OpenStudy (lncognlto):

Why is the integral of:

OpenStudy (lncognlto):

\[\int\limits_{x}^{y}(\frac{ 4 }{ {y}^{2}} - 1) = \left[ - \frac{ 4 }{ y } \right] -y\]

OpenStudy (lncognlto):

Sorry the -y at the end should be inside the parentheses.

OpenStudy (lncognlto):

the limits of y and x are not that important to this question right now, as I understand, so they are just place-fillers.

OpenStudy (lncognlto):

Why 4/y^2 becomes -4/y is what is confusing me.

OpenStudy (anonymous):

You know the integral of \[x^p\]is \[\frac{ 1 }{ p+1 }x^{p+1}\]Now think of \[\frac{ 4 }{ y^2}=4y^{-2}\rightarrow p=-2\]and apply the formula

OpenStudy (cggurumanjunath):